### v1

@@ 2.1.1 Analysis of fitting ability
Figure 1: Figure a illustrates the fitting capabilities and extrapolation properties on Eq.Nguyen-6 $sin(x^{2}+x)+sin(x)$ of various algorithms; Figure b depicts the change in algorithmic efficiency before and after the use of ada-$\alpha$; Figure c presents the correlation matrix among different variables in the Boston housing price prediction; Figure d demonstrates the schematic representation of housing price predictions using the variable ‘RM’. From the figure, it can be deduced that the housing prices are directly proportional to the variable ‘RM’, which is consistent with the findings presented in Figure c; Figure e displays the illustrative results of predicting housing prices using the variable ‘LSTAT’. Figure f showcases the schematic representation of predicting housing prices using both ‘RM’ and ‘LSTAT’. Figures g through n exhibit the prediction results of X-Net on univariate expressions; Figures o through v display the prediction outcomes of X-Net on multivariate expressions, where ‘-TV’ denotes true values and ‘-PRE’ represents predicted values.
2.1.2 Analysis of extrapolation ability
It is not enough to fit the given data well in the regression analysis. We also need the model to have good extrapolation ability. For example, we only have the data of $x$ in the range $[-2,2]$ and the corresponding value of $y$. We use this data as a training set to train the model. We need a model that not only works well in the range [-2,2], but also fits well outside the range [-2,2] (e.g., [-5,5]). We call this ability of the model extrapolation ability.
To test the extrapolation ability of the model, we sampled $x$ in the range of $[-2,2]$ and obtained the corresponding $y$ values. These data are then used to train each of the three algorithms X-Net, SVR, ANN, Kronecker NN, and Shinichi[6]. Finally, we sampled $x$ on $[-5,5]$ as a test set to test the three models and recorded the goodness-of-fit of the three algorithms in the interval $[-2,2]$ and outside the interval respectively, and the specific results are shown in Table 1. From the table, we can find that our algorithm is superior to SVR, ANN, Kronecker NN, and Shinichi[6] in extrapolation ability. The main reason for this is that the activation function of X-Net is more diverse and has more powerful representation ability. And the purpose of our algorithm is not only fitting but to find a mathematical expression that can well represent these data. And, the expression has good extrapolation, and its representation ability will not be significantly affected by the change in the $x$ value range. The simple diagram is shown in Fig. 1a
2.1.3 Diversity activation function representation capability analysis
To explore the powerful representation power of activation function diversity, we designed a unique test scheme: remove the weight parameter $W$ and bias term $B$ from a neural network, and try to fit the data in Table 1 purely by selecting and optimizing the activation function type of each neuron node, taking advantage of the nonlinear nature of these activation functions. This process only depends on the nonlinear expressive power of the activation function. From the last column of Table 1 ‘X-Net (No parameters)’, we can find that X-Net can make a good fit to the data and show good extrapolation even when no parameters are introduced. This experiment strongly demonstrates the powerful nonlinear representation ability of our neural network.
2.1.4 Comparison of network structure complexity
Artificial neural networks have a fixed network structure and a number of nodes, which can easily lead to redundancy in parameters and nodes, greatly slowing down the convergence efficiency of the neural network. In contrast, X-Net is more flexible, with a dynamically changing network structure and the number of nodes that can be adaptively adjusted based on the complexity of the problem. This greatly alleviates the problem of redundancy in nodes and parameters in ordinary neural networks. In order to test the complexity of the model obtained by four algorithms X-Net, ANN, Kronecker NN, and Shinichi[6] on the same task, we stop the training when $R^{2}=0.99$, and count the number of nodes and parameters used by the four networks at this time. The specific statistical results are shown in Table 2. From Table 2, it is evident that in fitting the Nguyen data set, the ANN necessitates approximately fourfold the node count in comparison to X-Net and demands almost a twentyfold increase in the quantity of parameters for optimization. The Kronecker NN algorithm demands roughly thrice the node count relative to X-Net and entails nearly a decuple increase in the number of parameters relative to X-Net. Shinichi[6] requires about three times as many nodes and five times as many parameters as X-Net.
2.2 Classification ability evaluation
The classification task stands as a quintessential sub-domain within the realm of machine learning. Consequently, gauging the classification performance of algorithms becomes imperative. To validate the classification efficacy of X-Net, we employed the Iris dataset and the MNIST handwritten digit recognition dataset as our experimental datasets. We juxtaposed the performance of X-Net with the ANN. On the MNIST dataset, we conducted two distinct experiments: initially, we employed PCA to dimensionally reduce the data, transmuting the images from a 28x28 dimension down to 6 dimensions, and subsequently classified using the reduced data. The secondary experiment involved direct classification without any dimension reduction. Table 2 meticulously delineates the performance comparison between X-Net and Others on both the Iris and MNIST datasets. The outcomes suggest that X-Net marginally outperforms the conventional ANN across all three classification tasks. Regarding network structural complexity, X-Net surpasses its counterpart. Taking the Iris dataset as a case in point, the neuron count of X-Net is merely half that of ANN, with its parameter magnitude being a mere tenth of that of ANN. On the MNIST dataset, In the experimentation employing PCA, the node count of X-Net is but a quarter of ANN, and its parameter magnitude stands at a mere $1.4\%$ (1/70) of ANN. In experiments without dimension reduction, the node count of X-Net amounts to two-thirds that of ANN, while the parameter count is a mere $2.9\%$ (3/100) of ANN. In summation, X-Net not only excels in classification tasks but also showcases a substantially reduced network structural complexity compared to the traditional ANN. Similarly, compared to Kronecker NN and Shinichi[6], X-Net also has significant advantages.
2.3 Used in scientific discovery
2.3.1 Modeling in economics (Boston house price forecast[9])
The Boston house price forecast data set is a very classic data set in the field of machine learning. Although it is simple, it can test the comprehensive performance of the algorithm well. Each sample in the Boston housing price data set is composed of 12 characteristic variables such as ’CRIM’, ’ZN’, and housing price MEDV. The thermal map of the correlation coefficient between each variable is shown in the figure1c. From the heat map 1c, we can find that the variable ’RM’ (the number of rooms per house) has the highest correlation coefficient with the MEDV(housing price), which is 0.70, showing a positive correlation with the housing price. In contrast, the variable ’LSTAT’ (how many people in the area are classified as low-income) has a minimum correlation coefficient of -0.74 with housing prices, showing a negative correlation with housing prices.
$MEDV=9.10RM-34.67(RM\geq 2)$ (2)
The predicted results are shown in Fig. 1d. From Formula 2, we can find that the final expression obtained by our algorithm can well reflect the positive correlation between RM and housing price.
We used LSTAT as input to predict housing prices using X-Net and finally got formula (3), from which we could see that the formula also well reflected the negative correlation between LSTAT and MEDV. Show in Fig 1e.
$MEDV=\frac{81.39}{(LSTAT+0.44)^{0.44}}-6.54(0\leq LSTAT\leq 100)$ (3)
Finally, we use RM and LSTAT as inputs to predict housing prices, and we get formula (4), And the predicted results are shown in Fig. 1f
$MEDV=\frac{15.6(0.48RM-13.21)}{2.36LSTAT-24.53}+9.97$ (4)
From Formula 4, we can clearly find that the housing price is directly proportional to the variable RM and inversely proportional to the variable LSTAT. That is the more rooms in the house, the more expensive the house, the more low-income people in the community, and the lower the price. The result of this prediction is completely in line with the facts. The above results prove that our algorithm can well reflect the relationship between each variable and the result even for multiple variables.
(a) Results of fitting the Airfoil-self-noise data (Eq.5)
(b) Figure a after sorting by $y_{pre}$
(c) Correlation coefficient matrix for Airfoil-self-noise
(d) Results of fitting the Airfoil-self-noise data(Eq.6)
(e) Figure d after sorting by $y_{pre}$
(f) Global temperature change prediction outcomes
(g) Solar energy power generation fitting result
(h) Figure g after sorting by $y_{pre}$
(i) Wind energy power generation fitting result
(j) Figure i after sorting by $y_{pre}$
(k) The correlation matrix in wind energy prediction
@@ 2.3.1 Modeling in economics (Boston house price forecast[9])
Figure 2: Figure a and Figure b show the fitting results of Formula 5 for Scaled sound pressure level; Figure c displays the correlation coefficient matrix for various variables of Airfoil-self-noise; Figure d and Figure e show the fitting results of Formula 6 for Scaled sound pressure level. Figure f displays the fitting results of the Formula 7 for global temperature changes; Figure g and Figure h show the prediction results of Formula 8 for solar power generation; Figure i and Figure j display the fitting results of Formula 8 for wind power generation data; Figure k represents the correlation coefficient matrix of the variables used in the wind power generation data.
2.3.2 Modeling in space science (Airfoil-self-noise)[10]
@@
The comparison between the predicted data and the real data for the two formulas above is shown in Figure 2. Figure 2a and Figure 2b demonstrate the fitting results of Equation 5. Note: To more clearly present the fitting outcomes of Equation 5, without altering the $R^{2}$ value, we sorted the data points in Figure a based on the predicted values $y_{pre}$ in ascending order, resulting in Figure 2b. From Figure 2b, it is evident that Equation 5 passes through the center of the data points, robustly fitting the data points. Figure 2d and Figure 2e depict the fitting results of Equation 6, with Figure 2e undergoing the same data processing as Figure 2b. From equation 5 above, we can observe that the predicted value y is directly proportional to $x_{1}$, $x_{2}$, $x_{3}$, and $x_{5}$, and inversely proportional to $x_{4}$. The situation reflected in the formula is consistent with the correlation coefficient heatmap 2c. From the graph, we can clearly see that only the fourth feature variable is directly proportional to the predicted result $Y$, while the other variables are inversely proportional to the result $Y$. Especially in equation 6, our algorithm only learned from four variables, $x_{1}$, $x_{3}$, $x_{4}$, and $x_{5}$, and missed variable $x_{2}$. Although it may not seem like a perfect result, we can see from the correlation heatmap that variable $x_{2}$ has a high correlation with $x_{5}$, with a correlation coefficient of 0.75, almost linearly related. Therefore, it is even wiser to retain $x_{5}$ and discard $x_{2}$ in the equation. This indirectly reflects the strong learning ability of our algorithm.
2.3.3 Modeling in environmental science (Prediction of Earth’s temperature change)[11].
@@
2.3.4 Modeling in energy science (Solar And Wind Power Generation Forecasting)[12]
$Y=x_{1}(0.00044x_{3}(-0.00003x_{1}x_{2}+0.31)+0.093)-0.2466$ (8)
The fitted curves of the sun power are shown in Figure 2g and 2h. Note: Figure h underwent the same data sorting process as Figure b.
The fitted curves of the wind power are shown in Figure 2i and 2j. Note: Figure j underwent the same data sorting process as Figure b. From Equation 9, we can see that X-Net has modeled the data well using only the first variable (wind speed) among the three variables. This is consistent with our understanding that wind power generation is mainly related to wind speed, which also aligns with reality. Furthermore, from the correlation coefficient heatmap 2k, we can clearly see that the correlation between $x_{1}$ (wind speed) and $y$ (power generation) is as high as 0.86.
2.4 Ada-$\alpha$ function performance analysis
In the process of neural network training, whether the step size is appropriate directly affects the speed of network convergence. If the step size is excessively large, the optimal solution may be skipped, If the step size is too small, the convergence rate may be extremely sluggish, and the optimal solution may not be obtained within a predetermined number of iterations. Therefore, to improve network efficiency, we design a step-size dynamic adaptive function Ada-$\alpha$. In order to test its performance, we selected four functions ’Nguyen-2’, ’Nguyen-6’, ’Nguyen-9’, and ’Nguyen-10’ from the Nguyen data set to test the Ada-$\alpha$. Using the Ada-$\alpha$ function, we execute each of the formulas ten times, record the amount of time required to fully recover the formula, and then calculate an average. Then do the same for the case without the Ada-$\alpha$ function. The specific bar graph is depicted in Fig. LABEL:fig9b. We can observe that when the Ada-$alpha$ function is implemented, the network’s convergence speed will increase dramatically.
3 Discussion
@@
We propose a new neural network model X-Net, and the network structure of X-Net is not fixed but dynamically adapts to changes in real-time under the guidance of gradients. The structure and parameter redundancy of traditional artificial neural networks is greatly improved. Specifically, we design the network structure as a tree and in order to improve the representation ability of the neural network, we extend the activation function of the neural network to the activation function library, which not only contains the traditional activation functions such as $[ReLU,sigmoid]$, but also includes the basic functions such as [+,,×,÷$[+,-,\times,\div,sin,cos,exp,sqrt,log]$. Next, we propose a new backpropagation mechanism, which can optimize not only the parameters of the network but also the activation function of the network nodes and the network structure. In particular, we take the output of each activation function as a variable $E$, differentiate the output of nodes of each layer through the chain rule, and then update it by the backpropagation algorithm. Finally, we select the activation function of nodes according to the updated $E$.
@@
X-Net outperforms state-of-the-art baselines in its ability to accurately fit the observed data. In particular, X-Net has better ”extrapolation” than traditional regression algorithms. In addition, in order to further test X-Net’s ability to aid scientific discovery, we tested X-Net’s performance in multiple scientific fields, including social science, environmental science, energy science, and space science. In the end, X-Net came up with a good mathematical formula for modeling data in various disciplines and accurately reflected the relationship between the variables $X$ and the predicted value $Y$ in the data. For observational data in all of these areas, X-Net was able to summarize mathematical formulas that were concise, analyzable, and consistent with objective facts.
Moreover, inspired by people descending the mountain and combined with the problems existing in the gradient descent method, we proposed a step-size dynamic adaptive function Ada-$\alpha$ that can obviously improve the efficiency of convergence and can be widely applied to other deep learning domains.
Although the X-Net input in the article is two-dimensional and the output is one-dimensional, in fact, the dimensions of the input and output of X-Net are not explicitly limited. For example, if we remove the primary node from a tree of symbolic expressions, the output of the network becomes two-dimensional. So X-Net can also handle problems where the output is multi-dimensional. As a result, X-Net is theoretically capable of doing much of what ANN can do.
In future studies, We are going to explore the application area of X-Net in a near step. For example, we can develop a new recurrent neural network based on X-Net. We can also develop a new convolutional neural network by replacing the last fully connected layer of the convolutional neural network, etc.
@@ 4 Methods
Similar to the ANN, our algorithm consists mostly of two components: (1) Initialize a Tree-like network structure and perform forward propagation. (2) Utilize the Alternating backpropagation technique to update the parameters and activation function alternately. (3) Repeat the above process until the maximum number of iterations is reached, or the goodness of fit $R^{2}$[13] reaches the set threshold and then stops the loop. The algorithm flow diagram for X-Net is shown in Fig. 3a.The detailed procedure of the algorithm is as follows: (i) The training data is obtained by sampling or observing. (ii) Obtain the data $x,y_{noise}$ to be input into the network, then use the binary tree of the formula $x^{2}+sin(x)$ as the initialization structure of the network, and note the initialization constants $W,B$ , where $W$ is initialized to $[1,1...1]$ and $B$ is initialized to $[0,0,...,0]$. Then, using forward propagation, the predicted value $y_{pre}$ is obtained, and MSE is calculated. (iii) Alternating Backpropagation, which involves taking partial derivatives with respect to constants and node outputs, If the condition $key\%ITE=0$ is not satisfied, then the constants are updated. If the condition is met, the outputs of nodes with the value $E$ are updated, followed by activation function updates via the Alternating Backpropagation mechanism. (iv) The previous procedure is repeated until a predetermined number of iterations or a predefined level of accuracy is obtained. Fig. 3b depicts the macro comparison between our technique and the ANN. Fig. 3c describes the forward propagation process of our algorithm in detail. In the figure, $E_{i}$ represents the output of the $i^{th}$ node, $w_{i}$ and $b_{i}$ represent the two constants of the $i^{th}$ node. The key distinction from ANN lies in the fact that we optimize not only the parameters W and b, but also treat the output E of each node as an optimizable variable, which undergoes alternate optimization with W and b. Subsequently, the activation function is updated based on the optimized E. A schematic of the dynamic adjustment of the network structure in the process of optimization is shown in Fig.3d. From the figure, we can see that the network structure and the neurons are adjusting in real time. We end up with the expression $y=sin(x^{2}+x)+sin(x)$.
@@
(b) Comparison between X-Net and ANN structures
@@
(d) Schematic diagram of structural changes during the training process
Figure 3: Figure a illustrates the algorithmic flowchart of X-Net; From Figure b, we observe a performance discrepancy between X-Net and ANN in the same regression task, with X-Net having a significantly lower network complexity compared to ANN; The Figure c describes the forward propagation process of our algorithm in detail; Figure d illustrates the dynamic changes in the network structure of X-Net during the training process.
@@ 4.1 Forward Propagation
Firstly, we randomly initialize a tree-like neural network, and the activation function of neurons is assigned according to the rules so that the neural network can realize normal forward propagation. For example, if the activation function is $[relu,sin]$ and so on, then the neuron only needs to have one input. But if the neuron is a binary activation function like [+,,×,÷][+,-$\times$,$\div$][ + , - , × , ÷ ], then the input must be two. Since the network structure is a binary tree, we can expand the network in the order of preorder traversal. We assume that $S=[s_{1},s_{2},...,s_{n}]$ is of the neural network . And there is a one-to-one correspondence between network structures and the preorder traversal of neural networks. $s_{i}$ in $S$ indicates that the neural network preorder traverses the $i^{th}$ Activation functions of neurons including unary activation functions $[ReLU,sigmoid,sin,cos,log,exp,sqrt,...]$, Binary activation function [+,,×,÷][+,-$\times$,$\div$][ + , - , × , ÷ ], and the leaf node symbols $(x_{1},x_{2},...,x_{n}]$. We start by randomly initializing a tree neural network, which in particular can be represented by a mathematical expression $x^{2}+sin(x)$. The network structure is shown in Fig. LABEL:fig3:env. For each node in the neural network, whether it is a leaf node or an intermediate node, we will have two constants, such as the node $sin(x)$ in the above network, for this node we have two parameters $d_{1}$ and $b_{1}$, namely its output is $d_{1}sin(x)+b_{1}$. After forward propagation, the final output of the neural network is $\hat{y}$. And then we define the loss function as the mean square error (MSE)[14] whose formula is as follows(10):
$Loss=\frac{1}{N}\sum_{i=0}^{N}{(y_{i}-\hat{y}_{i})}^{2}$ (10)
Where $N$ is the number of sampling points, $y_{i}$ is the true label value of the $i_{th}$ sample, and $\hat{y}_{i}$ is the predicted value of the $i^{th}$ sample. The objective is to continuously update the node symbols and parameters using the back-propagation technique so that the loss function value decreases over time.
(a) Unary functions to unary functions
(b) Unary functions to binary functions
(c) Unary functions to x
(d) Binary functions to unary functions
(e) Binary functions to binary functions
(f) Binary functions to x
(g) x to unary functions
(h) x to binary functions
Figure 4: This illustration depicts the evolutionary process among different neurons. This mechanism ensures that the neural network can still conduct forward propagation normally during its growth and evolution. The operator circled in red is the selected operator. Herein, the neuron types selected after differential gene expression are delineated by red boxes.

### v2

@@ 0.5.1 Modeling in space science (Airfoil-self-noise) (?)
Figure 2: Figure a and Figure b show the fitting results of Formula 2 for Scaled sound pressure level; Figure c displays the correlation coefficient matrix for various variables of Airfoil-self-noise; Figure d and Figure e show the fitting results of Formula 3 for Scaled sound pressure level. Figure f displays the fitting results of the Formula3 for global temperature changes; Figures g through n show the fitting results of X-Net on univariate benchmarks. Figures o through v display the prediction outcomes of X-Net on the multivariate benchmark, where ‘-TV’ denotes true values and ‘-PRE’ represents predicted values;
@@
The comparison between the predicted data and the real data for the two formulas above is shown in Figure 2. Figure 2a and Figure 2b demonstrate the fitting results of Equation 2. Note: To more clearly present the fitting outcomes of Equation 2, without altering the $R^{2}$ value, we sorted the data points in Figure a based on the predicted values $y_{pre}$ in ascending order, resulting in Figure 2b. From Figure 2b, it is evident that Equation 2 passes through the center of the data points, robustly fitting the data points. Figure 2d and Figure 2e depicts the fitting results of Equation 3, with Figure 2e undergoing the same data processing as Figure 2b. From equation 2 above, we can observe that the predicted value y is directly proportional to $x_{1}$, $x_{2}$, $x_{3}$, and $x_{5}$, and inversely proportional to $x_{4}$. The situation reflected in the formula is consistent with the correlation coefficient heatmap 2c. From the graph, we can clearly see that only the fourth feature variable is directly proportional to the predicted result $Y$, while the other variables are inversely proportional to the result $Y$. Especially in equation 3, our algorithm only learned from four variables, $x_{1}$, $x_{3}$, $x_{4}$, and $x_{5}$, and missed variable $x_{2}$. Although it may not seem like a perfect result, we can see from the correlation heatmap that variable $x_{2}$ has a high correlation with $x_{5}$, with a correlation coefficient of 0.75, almost linearly related. Therefore, it is even wiser to retain $x_{5}$ and discard $x_{2}$ in the equation. This indirectly reflects the strong learning ability of our algorithm.
0.5.2 Modeling in environmental science (Prediction of Earth’s temperature change) (?).
@@
Discussion
In this study, we conduct an in-depth exploration of next-generation neural networks and introduce a novel neural network model named X-Net. Theoretically, this model allows the use of any differentiable function as its activation function, which is not static but dynamically learned under the guidance of gradient information. Furthermore, the network structure of X-Net is dynamically adaptable, and capable of self-adjusting at the neuronal level in real-time, including both growth and reduction, to better adapt to specific tasks while minimizing the issues of parameter redundancy and insufficiency. Notably, when dealing with relatively simple tasks, X-Net can derive a simplified and interpretable mathematical formula, which is particularly beneficial for scientific research.
@@ Discussion
Specifically, we design the network structure as a tree and in order to improve the representation ability of the X-Net, we extend the activation function of the neural network to the activation function library, which not only contains the traditional activation functions such as $[ReLU,sigmoid...]$ but also includes the basic functions such as [+,,×,÷$[+,-,\times,\div,sin,cos,exp,sqrt,\\ log...]$. Next, we propose an alternating backpropagation mechanism, which can optimize not only the parameters of the network but also the activation function of the network nodes and the network structure. In particular, we take the output of each activation function as a variable $E$, differentiate the output of nodes of each layer through the chain rule, and then update it by the backpropagation algorithm. Finally, we select the activation function of nodes according to the updated $E$.
@@
X-Net achieves comparable performance to MLP for both classification and regression. However, in terms of network structure complexity, X-Net is far less than MLP. In addition, to further test X-Net’s ability to aid scientific discovery, we tested X-Net in multiple scientific fields, including social science, environmental science, energy science, and space science. In the end, X-Net came up with a concise, analyzable mathematical formula for modeling data in various disciplines and accurately reflected the relationship between the variables $X$ and the predicted value $Y$ in the data.
X-Net has great application potential. For example, we can use it to do algorithmic distillation, to distill a complex network into a simple X-Net. Or we could try to replace the fully connected layers in Transformers, GPT, or other deep learning networks and drastically reduce the model size. It can speed up the inference efficiency and reduce the deployment cost of the model. We can also use it to solve partial differential equations(PDE) so that it is possible for us to obtain an analytical solution to the partial differential equation. We can also apply X-Net to healthcare, finance, and other fields with high interpretability requirements to open up the MLP black box system. In summary, we can theoretically use X-Net to do a lot of things that MLP can do.
X-Net also has a number of drawbacks, such as training is also relatively time-consuming. There are even occasional cases where the training fails to converge.
In summary, X-Net provides a new perspective and opens a new possibility for the study of next-generation neural networks. We sincerely hope that X-Net can be a little inspiring to the following researchers.
@@ 1 Methods
As shown in Fig. 3a, X-Net is composed of several steps: (1) Feed the training data $\{x,y_{\rm noise}\}$ which are obtained from the sampling of mathematical expressions or observations of real-world to the network; (2) Initial the tree-shape network structure and constants, denote the constants as $W$ (weights) and $B$ (bias); (3) Perform a forward propagation to obtain the predicted value $y_{\rm pre}$; (4) Compute the loss by loss function; (5) Implement the alternating backpropagation algorithm to alternately update parameters $W,B$ as well as the activation function through updating the node output value $E$; (6) Repeat steps (3)-(5) until X-Net exceeds the expected $R^{2}$  (?) or reaches the maximum epoch.
X-Net is a tree-shape network and we update the network through the backpropagation algorithm similar to MLP. The distinctions lie in the backpropagation procedure, in which we add an extra step to substitute proper activation functions by the guide of gradients. Fig. 3b depicts the structure of X-Net and MLP, respectively. Compared with MLP, the connection of X-Net is sparse and the activation functions are diversiform. Fig. 3c shows the forward propagation of X-Net, the $E_{i}$ represents the output of $i^{\rm th}$ node, and $w_{i}$ and $b_{i}$ are the constants. X-Net optimizes the parameters $W$ and $B$ and the activation functions, which is the key distinction with MLP. Both the network structure and the neuron types are adjusted during the learning process.
@@
(b) Comparison between X-Net and MLP structures
@@
Figure 3: Figure a illustrates the algorithmic flowchart of X-Net; From Figure b, we observe a performance discrepancy between X-Net and MLP in the same regression task, with X-Net having a significantly lower network complexity compared to MLP; Figure c describes the forward propagation process of our algorithm in detail;
@@ 1.1 Forward Propagation
We use the pre-order traversal of the X-Net to code since it’s a tree-shape network and with a unique mapping between the binary tree and pre-order traversal. Assume the pre-order traversal of a X-Net is $S=[s_{1},s_{2},...,s_{m}]$, where $m$ is the number of neuron or number of node, $s_{i}$ indicates the $i^{\rm th}$ neuron. The activation function of each $s_{i}$ is selected from a library $\{ReLu,sigmoid,sin,cos,log,exp,sqrt,...,+,-,\times,\div,x_{1},x_{2},...\}$ that contains unary, binary, and variables. X-Net is initiated according to the arity of the activation function of each neuron: if the activation function requires two inputs, such as +,,×,÷+,-$\times$,$\div$+ , - , × , ÷, then the neuron has two inputs; otherwise, the neuron has one input. Each neuron will have two distinct constant values $w_{i}$ and $b_{i}$. The forward propagation is conducted from left to right (as shown in Fig. 3c), or alternately, from leaf to root for the binary tree. The root node outputs the predicted value $\hat{y}$. Finally, the loss function is calculated according to different types of tasks.